A balloon is inflated to a volume of 1.250 L with dry hydrogen gas, at 28.0 °C. The balloon is then
cooled, and its volume drops to 964 mL. If the pressure is unchanged, what is the hydrogen gas temperature (in °C) in the cooled balloon? Assume that no gas has been added or removed.
1
Expert's answer
2012-05-22T12:08:57-0400
PV=nRT
n=const
(P1V1)/T1=(P2V2)/T2
P=const
V1/T1=V2/T2
1.250/(273+28)=0.964/x

x=232.3
K= -40.7 C